9. A solution of ethyl alcohol (C2H5OH) in water has a concentration of 20.56% v/v. What volume of C2H5OH is present in 255 mL of solution?

10. What mass of KCl is present in 475 mL of a 1.09% m/v aqueous solution?

11. What is the molarity of 0.500 L of a potassium chromate solution containing 0.0650 mol of K2CrO4?

12. What is the molarity of 4.50 L of a solution containing 0.206 mol of urea [(NH2)2CO]?
Answer #1

9)

Given:

(v/v)% = 20.56 %

Volume of solution = 255 mL

Use:

(v/v)% = volume of C2H5OH * 100 / volume of solution

20.56 = volume of C2H5OH * 100 / 255

volume of C2H5OH = 52.4 mL

Answer: 52.4 mL

10)

Given:

(M/v)% = 1.09 %

Volume of solution = 475 mL

Use:

(M/v)% = mass of KCl * 100 / Volume of solution

1.09 = mass of KCl * 100 / 475

mass of KCl = 5.18 g

Answer: 5.18 g

Only 1 question at a time please